OpenStudy (anonymous):

How many solutions does this equatiom have? 0=-4b+4b

OpenStudy (anonymous):

Infinite. Because no matter what value b will be, it will always be true.

OpenStudy (anonymous):

can u explain how u got that answer

OpenStudy (anonymous):

\[r+11+8r=29\]

OpenStudy (anonymous):

Take 0=-4b+4b Substitute b for any number you can think of, will it still be equal to 0?

OpenStudy (anonymous):

I seriously dont understand wish someone would teach me!!!!!

OpenStudy (anonymous):

What number do you want b to be? pick anything.

OpenStudy (anonymous):

5

OpenStudy (anonymous):

\[0=-4\times5+4\times5\] \[0 = -20+20\] \[0 =0\]

OpenStudy (anonymous):

ok so anyway u can explain in detail how u come up with that i am starting to catch on

OpenStudy (anonymous):

Do you know how to solve simple equations?

OpenStudy (anonymous):

x+1=2

OpenStudy (anonymous):

Solve that one for x.

OpenStudy (anonymous):

how about 6r-1=6r=11 these i honestly dont understand how to solve

OpenStudy (anonymous):

Trust me I want to help you understand the basics, from there you'll solve any equation.

OpenStudy (anonymous):

please

OpenStudy (anonymous):

So what would be the solution to x+1=2

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

How many solutions does the equation have?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Solution means how many x=number are there.

OpenStudy (anonymous):

In this case x=1

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

Meaning there's only one solution. If we take 1 and plug it into the original equation we will get this: x+1=2 (1)+1=2 2=2 A true statement.

OpenStudy (anonymous):

1 is the only number in the world you can replace x with to get a true statement

OpenStudy (anonymous):

So it has only one solution, x=1.

OpenStudy (anonymous):

okay, i need to figure out what r equals

OpenStudy (anonymous):

In here? r+11+8r=29

OpenStudy (anonymous):

Or did you have a different question?

OpenStudy (anonymous):

okay, 6r-1+6r=11 i need to figure out what r equals

OpenStudy (anonymous):

Do you know how to combine like terms?

OpenStudy (anonymous):

kind of

OpenStudy (anonymous):

12r-1=11

OpenStudy (anonymous):

Add +1 to both sides of the equation.

OpenStudy (anonymous):

12r=12

OpenStudy (anonymous):

divide both sides by 12

OpenStudy (anonymous):

r=1

OpenStudy (anonymous):

If you substitute 1 for r in the original equation you will get 11=11

OpenStudy (anonymous):

Meaning that there is only one solution to that equation.

OpenStudy (anonymous):

okay, thanks